b) Evren wants to go into the donut business. For $500 per month he can rent a bakery complete with all the equipment he needs to make a dozen different kinds of donuts (K-1, r-500). He must pay unionized donut bakers a monthly salary of $400 each. He projects his monthly production function to be Q 5KL, where Q is tons of donuts. i) With the current level of capital, what is the marginal product of labor (MPL)? Is the 3 marks) i) If Evren wishes to make 25 tons of donuts, how many bakers are required given the 4 marks) marginal product diminishing? current level of capital? How much will it cost to produce this (total cost)? ii) Derive Evrens short-run cost function with K- (3 marks)c) Draw a graph showing the average total cost, average variable cost, and marginal cost curves for a typical firm. Draw in THREE (3) prices that result in the firm making positive profits, (5 marks) d) Suppose a firm has the following total cost function: TC 100 +4qf. What is the minimum breaking even, and making negative profits that are less than fixed costs price necessary for the firm to earn profit? Below what price will the firm shut down in the (5 marks) e) If each competitive firm in an industry has the short-run cost function C 505q q, and short run? the market price is $35, what is the profit-maximizing output level for each firm? What is the (5 marks) f) Assume a monopolist faces a market demand curve P 100 - 2Q and has the short-run total total revenue? Set MC-MR and solve cost function C 640 + 20Q i)What is the profit-maximizing level of output? What is the profits? Graph the marginal revenue, marginal cost, and demand curves, and show the area that (7 Marks) In Question f (i), what would price and output be if the firm priced at socially represents deadweight loss on the graph ii) efficient (competitive) levels? What is the magnitude of the deadweight loss caused by monopoly pricing? (3 Marks)

Please explain in details with step-by-step solution, Thank you very much
Answer #1

; ) k-19(7 50 o , l@ w= 4DO, &= 5kl dhuu Mf being Constaut, S0 3 &not diminishiug A OL Now total Cost Cょ U) Lat qKF (400x5) +Foo/1 8000 +500 = 3500 t 500 = 8o.t 500answer c)

TT 9 Vc making 3ovo T, it be eaka 8-9, not eUem V0%iable cost ae cowerede) e-50 t5112 p-35, thus MR-P-3s , 35= 5+89/ 35X1555 Bo Total Revenue - Pa Aanswering only first four parts as per HomeworkLib policy